What resistance must be placed in parallel with a 151 Ω resistor to make the equivalent resistance 128 Ω?
1
Expert's answer
2013-02-07T08:15:26-0500
Let's denote the unknown resistance by R. Then according to the Ohm's law,

1/151[Ω] + 1/R[Ω] = 1/128[Ω] ==>

R[Ω] = 1/(1/128[Ω] - 1/151[Ω]) = 19328/23 Ω (≈ 840.34 Ω).

So, 19328/23 Ω resistor must be placed in parallel with a 151 Ω resistor to make the equivalent resistance 128 Ω.